a car travelled 450 miles in 7.5 hours. a traffic accident cut the car's driving speed in half for 1.5 hours. what was the car's speed for the remainder of the trip?
Answer by fractalier(2101)

Remember that distance = rate * time, or D = RT.
So if the car travels 450 miles in 7.5 hours, its rate is
R = D/T = 450/7.5 = 60 mph
If that speed is cut in half, the new rate is 30 mph.